# Black holes immersed in dark matter: energy condition and sound speed

### v1

@@ VI Discussion and conclusion
Figure 10: In the first row we show the behavior of the dominant energy condition and the $c_{st}^{2}$. At the bottom, we show $|c_{st}^{2}|$ and the derivative of density with respect to the radius. The connection between the vanishing $\rho^{\prime}$ and the diverging sound speed is present here also. The profile is qualitatively similar to all the other profiles. Therefore, all the findings in the current work are qualitatively valid for the NFW profile.
Figure 11: In the first row we plot $c_{st}^{2}$ and we zoom in the diverging region. For all the profiles, below a certain radius $c_{st}^{2}$ becomes negative. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. Note, that the region of divergence is shifted to a larger radius for $(1-4/r)$ scaling compared to the $(1-2/r)$ scaling.
@@
In Ref. [23, 24] it was demonstrated that the DM density near BH vanishes. Based on this, previous works had scaled the density profile to make the density vanish in the near zone. We followed the same for the latter half of our work. We took these modified density profiles. We found before imposing the sound speed condition that these profiles have more pathologies. Not only does the dominant energy condition and the tangential sound speed become superluminal but also the square of sound speeds becomes negative in the near region. By plotting the radial derivative of density we demonstrated these divergences and the negative values are connected to the flattening and decreasing nature of the density profile. Once we put the physical sound speed condition we find the pathologies are resolved and rather than a flattened and decaying density structure the density starts to grow. This points towards the possibility that the physical sound speed under the current modeling does not allow the density to get flattened or demonstrate a vanishing structure. It therefore contradicts the notion that the density should get flattened near BH. It remains to be seen how the possible contradiction with Ref. [23, 24] be addressed. It can have a connection with the “Cusp-Core” problem. It is also important to investigate in detail if the current modeling is appropriate for the DM energy-momentum tensor. It requires further detailed investigation.
It is also noteworthy that we have not assumed the solution to be a BH apriori. Along with the energy-momentum tensor model, we have only assumed the metric to be spherically symmetric that is asymptotically flat. From the behavior of $g_{rr}$ we know that it diverges at $r\gtrsim 2$ since $m(r)>1$. From the solution of $g_{tt}$, we indeed find that $g_{tt}\to 0$ as $r\to 2$. However, this does not imply apriori that the points where $g_{tt}=0$ and $g^{rr}=0$ are the same point. Hence, it requires further investigations focusing on the near zone to establish if it is indeed a BH or a wormhole or shows some other structure. This needs more investigation.
@@ Appendix A NFW
Appendix B Sadeghian-Ferrer-Will type profile
@@
In the main text along with the actual density profiles, we also showed the results of scaled density profiles. However, the scaling was done in a manner so that the density vanishes at the horizon, i.e. $r=2$. In Ref. [24] they found the density to vanish rather at $r=4$. For this purpose here we show the result for such kind of vanishing density. Here we take $\rho_{Hern/NFW}\to(1-4/r)\rho_{Hern/NFW}$. From the study in the main text, we showed that the vanishing of density implies there is a maximum in the density and the density decreases afterward. We demonstrated that this is the region where the sound speed behaves unphysically, precisely because of the density structure. Therefore having a different scaling will not change the qualitative features. The region of divergences will just get shifted.
@@
In the first row of Fig. 11 we plot $c_{st}^{2}$ and we zoom in on the diverging region. For all the profiles, below a certain radius, $c_{st}^{2}$ becomes negative even for the current scaling. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. These features are exactly similar qualitatively to the $(1-2/r)$ scaling we discussed in the main text. In the current scaling, the density vanishes everywhere $r<4$, unlike the previous case where it vanishes at the horizon. As a result, in the current case, the region of divergence is shifted to a larger radius compared to the $(1-2/r)$ scaling. This bolsters the point that flattening and vanishing characteristics are the reasons behind the unphysical behavior of sound speed.
@@
Interestingly for this scaling, in the region with the radius smaller than that of the divergence points, there exists a patch where sound speed can be physical. In the subplot of Fig. 11 this region is shown. Below the divergence point, the sound speed becomes negative similar to the other scaling. However, later it starts to grow and around $r\sim 5.5$, it reaches zero. From $\sim 5.5$ to $r=4$ it becomes positive and subluminal. We are not showing the region below $r=4$, as the density vanishes in that region. This result implies that depending on the scaling, below the divergence points physical sound speed can exist in certain patches. However, it can not resolve the divergences and the negative sound speed. Therefore the flatness of the density is bound to create these pathologies. However, depending on how and where the density gets flattened, physical sound speed patches can form near the BHs.

### v2

@@ VI Discussion and conclusion
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 10: The results for the NFW profile are shown here. In the first row, we show the behavior of the dominant energy condition and the $c_{st}^{2}$. At the bottom, we show $|c_{st}^{2}|$ and the derivative of density with respect to the radius. The connection between the vanishing $\rho^{\prime}$ and the diverging sound speed is present here also. The profile is qualitatively similar to all the other profiles. Therefore, all the findings in the previous part of the current work are qualitatively valid for the NFW profile.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 11: Here the results for SFW-type scaling $(1-4/r)$ are shown. In the first row, we plot $c_{st}^{2}$ and we zoom in the diverging region. For all the profiles, below a certain radius $c_{st}^{2}$ becomes negative. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. Note, that the region of divergence is shifted to a larger radius for $(1-4/r)$ scaling compared to the $(1-2/r)$ scaling.
@@
In Ref. [23, 24] it was demonstrated that the DM density near BH vanishes. Based on this, previous works had scaled the density profile to make the density vanish in the near zone. We followed the same for the latter half of our work. We took these modified density profiles. We found before imposing the sound speed condition that these profiles have more pathologies. Not only does the dominant energy condition and the tangential sound speed become superluminal but also the square of sound speeds becomes diverging and negative in the near region. By plotting the radial derivative of density we demonstrated these divergences and the negative values are connected to the maxima of the density profile. Once we put the physical sound speed condition we find the pathologies are resolved and rather than a maxima and decaying density structure the density starts to grow. This points towards the possibility that the physical sound speed under the current modeling does not allow the density to decrease or demonstrate a vanishing structure. It therefore contradicts the notion that the density should reach maxima and then vanish near BH. It remains to be seen how the possible contradiction with Ref. [23, 24] be addressed. It can have a connection with the “Cusp-Core” problem. It is also important to investigate in detail if the current modeling is appropriate for the DM energy-momentum tensor. It requires further detailed investigation.
Ref. [23, 24] considered the BH to be growing adiabatically inside the DM halo, which resulted in the density maxima near the horizon. On the other hand, in Ref. [53, 54] it has been shown that the regularity of $T^{\mu\nu}T_{\mu\nu}$ near horizon for a time-dependent spherically symmetric metric, i.e. $(g_{tt},g_{rr})$, requires non-zero $T_{rt}$ component. Since the vanishing density near the horizon comes from the calculations that assume adiabatic time evolution of the BH, it is important to investigate whether modeling such a DM profile as a diagonal energy-momentum tensor is appropriate. The origin of the violation of the energy condition and the ill-behaved sound speed could also be the diagonal form of the energy-momentum tensor, at least partially. Hence, it boils down to investigating the accuracy of the assumption that the DM profile is a static structure and the resulting metric is time independent too. These aspects must be investigated in the future.
Notably, we have not assumed the solution to be a BH apriori. Along with the energy-momentum tensor model, we have only assumed the metric to be spherically symmetric and asymptotically flat. From the behavior of $g_{rr}$ we know that it diverges at $r\gtrsim 2$ since $m(r)>1$. From the solution of $g_{tt}$, we indeed find that $g_{tt}\to 0$ as $r\to 2$. However, this does not imply apriori that the points where $g_{tt}=0$ and $g^{rr}=0$ are the same point. Hence, it requires further investigations focusing on the near zone to establish if it is indeed a BH or a wormhole or shows some other structure. This needs more investigation.
@@ Appendix A NFW
Appendix B Sadeghian-Ferrer-Will (SFW) type profile
In the main text along with the actual density profiles, we also showed the results of scaled density profiles. However, the scaling was done in a manner so that the density vanishes at the horizon, i.e. $r=2$. Ref. [24] studied what will be the density profile of a halo if a black hole is grown adiabatically within a preexisting halo. They found the density to vanish rather at $r=4$. The disappearance of particle density at $r=4$ is understood in terms of stable circular orbits. This value corresponds to the radius of the unstable circular orbit within the Schwarzschild geometry for a particle on the verge of escaping, characterized by energy per unit mass $E=1$ and having an angular momentum per unit mass $L=4$. A particle with $L\geq 4$ and $E\leq 1$ has an inner turning point at $r\geq 4$. Consequently, any particle managing to approach $r=4$ inevitably falls into the black hole. Therefore, the particle density vanishes at $r\leq 4$. For further details see [24].
@@ Appendix B Sadeghian-Ferrer-Will (SFW) type profile
For this reason, here we show the result for such kind of vanishing density. Here we take $\rho_{Hern/NFW}\to(1-4/r)\rho_{Hern/NFW}$. From the study in the main text, we showed that the vanishing of density implies there is a maximum in the density and the density decreases afterward. We demonstrated that this is the region where the sound speed behaves unphysically, precisely because of the density structure. Therefore having a different scaling will not change the qualitative features. The region of divergences will just get shifted.
@@
In the first row of Fig. 11 we plot $c_{st}^{2}$ and we zoom in on the diverging region. For all the profiles, below a certain radius, $c_{st}^{2}$ becomes negative even for the current scaling. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. These features are exactly similar qualitatively to the $(1-2/r)$ scaling we discussed in the main text. In the current scaling, the density vanishes everywhere $r<4$, unlike the previous case where it vanishes at the horizon. As a result, in the current case, the region of divergence is shifted to a larger radius compared to the $(1-2/r)$ scaling. This bolsters the point that the maxima of density is the reason behind the unphysical behavior of sound speed.
@@
Interestingly for this scaling, in the region with the radius smaller than that of the divergence points, there exists a patch where sound speed can be physical. In the subplot of Fig. 11 this region is shown. Below the divergence point, the sound speed becomes negative similar to the other scaling. However, later it starts to grow and around $r\sim 5.5$, it reaches zero. From $\sim 5.5$ to $r=4$ it becomes positive and subluminal. We are not showing the region below $r=4$, as the density vanishes in that region. This result implies that depending on the scaling, below the divergence points physical sound speed can exist in certain patches. However, it can not resolve the divergences and the negative sound speed. Therefore the maxima of the density is bound to create these pathologies. However, depending on how and where the density reaches maxima, physical sound speed patches can form near the BHs.